Calculate the ?G

Calculate the ?G
Answer #1

2 H2S + 3 O2 = 2SO2 + 2H2O

Delta Go recation = summation [delta Go formation of products] - summation [ delta Go formation of reactants]

for products = 2* -300.1 + 2*-228.6 = -1057.4 kJ/mol

for reactants = 2* -33.4 = -66.8 kJ/mol

delta G o reaction = -1057.4 + 66.8 = -990.6 kJ/mol